OpenStudy (anonymous):

When 3.651 grams of a hydrocarbon, CxHy, were burned in a combustion analysis apparatus, 11.46 grams of CO2 and 4.691 grams of H2O were produced. In a separate experiment, the molar mass of the compound was found to be 70.13 g/mol. Determine the empirical formula and the molecular formula of the hydrocarbon.

OpenStudy (anonymous):

Get the moles of C and H from the moles of CO2 and H20. Their ratio will give you x/y. You can get added information from the weight of O in the compounds given off. Then (x)(12) + (y) (1) = 70.13 will give you a second equation from these molecular weights.

OpenStudy (anonymous):

you add the the grams of C AND H and the given number of 3.561 to get the grams of O right

OpenStudy (anonymous):

To get the amount of O in the CO2, used the molecular weight (12 +16 + 16)= 44 to divide into the weight of CO2 found, 11.46g, to indicate there were 11.46/44 = 0.25 moles of CO2, thus 0.25 moles of C. Do the same with H2O (m.w. = 2 + 16=18) and find the # moles of H2O, which will mean there were twice this many moles of H and one times this many moles of O. You now have number of C and H moles, giving you x/y. You also have number of O moles, might be useful. [I have not done the whole problem and hope to go to bed soon,] Good luck.